# Simple Techniques for Enhancing Sentence Embeddings in Generative Language Models

### v1

@@ Abstract
The advent of generative Pre-trained Language Models (PLMs), exemplified by GPT [1], OPT [2], and LLaMA [3], marks a significant milestone in the field of Natural Language Processing (NLP). Boasting immense parameter sizes and extensive pre-training corpora, these models have exhibited unparalleled proficiency in multi-task processing and zero-shot reasoning. As a result, there is a growing interest in leveraging these models for a wide array of downstream tasks, including the derivation of sentence embeddings, without the necessity for gradient updates.
@@
Over the past few years, research in this direction has predominantly revolved around discriminative models like BERT [4] and RoBERTa [5]. Nonetheless, the semantic space anisotropy inherent in these models [6] often mandates fine-tuning with considerable datasets to produce high-quality sentence embeddings. Among various fine-tuning approaches, contrastive learning, which aims to bring similar samples closer while distancing disparate ones, is most commonly employed. However, in order to avert model collapse and ensure sufficient reference items for sentence vector calculation, contrastive learning methods necessitate large training batch sizes, thereby consuming massive computational resources.
@@
In response to these challenges, some cutting-edge efforts have shifted their focus to the recently developed generative models, hoping to leverage their advanced text comprehension abilities to directly encode input sentences without additional backward propagation. Given the disparity between sentence representation and language modeling, this endeavor demands ingeniously designed prompts. The pioneering work, PromptEOL [7], find that by introducing an Explicit One word Limitation (EOL) into the prompt and constructing a template akin to ‘This sentence: “[X]” means in one word: “’, it can significantly enhance the embedding obtained from generative models. Subsequent studies [8] [9] have followed this practice while also combining EOL with contrastive learning for fine-tuning Large Language Models (LLMs). Nevertheless, even with efficient fine-tuning strategies like QLoRA [10], training a neural network with a 7B-scale LLM as its backbone still consumes more GPU memory than fully fine-tuning a 110M-scale BERT. Moreover, we elucidate in subsequent sections that EOL chiefly benefits raw embeddings from generative models and is not requisite for discriminative models or fine-tuning generative models.
To reconcile the demand for high-quality sentence embeddings with the need for computational resource conservation, direct inference using generative models presents a compelling alternative. MetaEOL [11], a recent innovation building upon PromptEOL, employs ChatGPT-4 to craft diverse templates for selected NLP tasks and subsequently derives the corresponding embeddings from LLMs. By averaging these embeddings, it procures the final sentence representation. Although this approach achieves certain improvements, both the selection of specific downstream tasks and the prompt design for ChatGPT require a significant investment of human effort. Moreover, encoding the same sentence multiple times significantly impairs computational efficiency.
@@
Contrastingly, this paper designs two simple yet effective prompt engineering methods: Pretended Chain of Thought (CoT) and Knowledge Enhancement, to ameliorate the sentence representation of generative models under direct inference conditions. These techniques involve appending a fixed prefix before the EOL to harness the PLM’s contextual learning ability to its fullest potential. Specifically, Pretended CoT, inspired by the design of Zero-shot CoT [12], mainly serves an “inspirational” role since we do not need the model to output intermediate reasoning steps. Knowledge Enhancement, on the other hand, provides explicit guidance to the model by conveying human experience in text summarization through prompts. Our comprehensive experimental outcomes affirm that Pretended CoT and Knowledge Enhancement not only optimize memory efficiency (refer to Table 1) but also significantly elevate the performance of raw embeddings produced by generative models across multiple Semantic Textual Similarity (STS) benchmarks.
@@
Our investigations reveal that the widely applied Explicit One-word Limitation (EOL) is actually a sentence embedding derivation technique for direct inference scenarios with generative PLMs. It does not always yield optimal performance in discriminative models or fine-tuning settings for generative models. This finding opens new avenues for future research on manual template design.
This paper proposes two concise and potent prompting engineering methods, Pretended CoT and Knowledge Enhancement, that markedly enhance the quality of embeddings generated by generative models on top of the original prompt.
@@ 2 Related Work
The exploration of sentence embeddings utilizing PLMs constitutes a vibrant area of research within Natural Language Processing and Computational Linguistics. This body of work diverges into three primary streams: supervised fine-tuning, unsupervised fine-tuning, and direct inference. Each paradigm offers unique insights while presenting distinct challenges, as delineated below.
@@
Supervised Fine-tuning: This strand of research leverages high-quality annotated corpora to enhance models’ encoding capabilities. A landmark in this domain, Sentence-BERT [13] designs a Siamese neural network architecture with shared parameters and updates model parameters via the NLI dataset [14] [15]. Subsequent to this, SimCSE [16] filters the NLI dataset and organizes it into triplets to adjust the sentence vector distribution of BERT and RoBERTa through contrastive learning. Further studies like PromptEOL [7], AnglE [8], and DeeLM [9] extend this methodology to generative PLMs such as OPT [2], LLaMA [3], and LLaMA2 [17]. While supervised learning methods consistently outperform across a spectrum of benchmarks, their dependency on labeled datasets circumscribes their applicability in a broader context.
@@
Unsupervised Fine-tuning: Focusing on the potential of models trained with unlabelled corpora, unsupervised fine-tuning has emerged as a fertile ground for innovation. This realm is particularly enriched by contrastive learning methods and their integration with data augmentation techniques. ESimCSE [18], for example, advances SimCSE by introducing random word repetition in the input text to construct positives for original sentences. PromptBERT [19], in contrast, utilizes different manual templates to create positive instances. Additionally, SSCL [20] demonstrates the efficacy of using intermediate layer outputs from PLMs as hard negatives in the InfoNCE Loss [21] framework. CoT-BERT [22], representing the current state-of-the-art, nearly matches the performance of supervised approaches across multiple STS benchmarks by relying solely on the intrinsic capabilities of PLMs. Despite these encouraging advancements, unsupervised fine-tuning methods do not inherently resent advantages in memory efficiency, given their reliance on large batch sizes to facilitate effective contrastive learning.
@@
Direct Inference: Methods that enable efficient inference by eschewing parameter updates and directly utilizing raw PLM embeddings as sentence representations are increasingly capturing the interest of the academic community. Among these, PromptBERT [19] showcases that meticulously designed manual templates tailored to mimic the Masked Language Modeling (MLM) task can exploit BERT’s original output vectors for commendable performance. Building on this insight, PromptEOL [7], which targets generative PLMs, advocates for explicitly prompting the model to condense the original sentence into a single word, markedly tapping into the latent potential of generative models and surpassing the direct inference results of PromptBERT. Considering the multifaceted nature of refining textual semantics, MetaEOL [11] utilizes ChatGPT-4 to craft diverse templates, feeding them into the PLM and extracting the sentence representation through average pooling. Although the premise of this method is theoretically sound, its operational requirement to encode the same sentence multiple times with distinct templates diverges from the original intention of offline computation. Specifically, for a given query , MetaEOL needs to concatenate with unique templates, inputting them into the PLM for computationally intensive encoding prior to deriving the corresponding embedding for similarity assessments. The theoretical time consumption is times that of the conventional solution, severely impeding the efficiency of large-scale information retrieval.
@@ 3 Methodology
In this section, we first elucidate, through pilot experiments in Sections 3.1 and 3.2, that the primary function of the Explicit One-word Limitation (EOL) is to enhance the direct inference effects of generative PLMs. Subsequently, in Section 3.3, we provide detailed descriptions of two innovative prompting engineering techniques proposed in this paper, Pretended Chain of Thought (CoT) and Knowledge Enhancement, for further elevating the semantic expressiveness of PLMs’ raw embeddings.
@@ 3.1 EOL in the Context of Discriminative Models
The design philosophy of EOL stems from an analysis of inference outcomes when employing an unmodified BERT model on the STS-B development set. This analysis contrasts two highly similar manual templates, distinguished solely by the presence or absence of a terminal period:
@@
Expanding upon this premise, our study embarks on a more granular investigation. We commence with a series of comparative experiments centered on BERT, examining the impact of the number of [MASK] tokens and the selection of terminal characters in templates. Adopting the identical experimental setup as PromptEOL, we utilize ‘This sentence: “[X]” means [MASK].’ as our baseline template and initialize BERT from its original pre-trained state with its performance measured on the STS-B development set [32] via Spearman correlation coefficient. The corresponding results are presented in Table 2.
| [MASK] | EOS | Aggregation | STS-B dev |
| 1 | None | 27.21 | |
| 1 | [SEP] | 32.54 | |
| 1 | . | 73.61 | |
| 2 | . | average | 73.94 |
| 3 | . | average | 74.50 |
| 4 | . | average | 71.83 |
| 2 | . | concatenate | 73.40 |
| 3 | . | concatenate | 73.49 |
| 4 | . | concatenate | 69.61 |
| 1 | 77.59 | ||
| 2 | average | 76.55 | |
| 3 | average | 75.99 | |
| 4 | average | 73.45 | |
| 2 | concatenate | 76.36 | |
| 3 | concatenate | 75.02 | |
| 4 | concatenate | 71.45 | |
| 1 | 75.32 | ||
| 2 | average | 75.19 | |
| 3 | average | 75.37 | |
| 4 | average | 73.62 | |
| 2 | concatenate | 74.89 | |
| 3 | concatenate | 74.62 | |
| 4 | concatenate | 72.35 |
In Table 2, the column labeled “[MASK]” denotes the number of [MASK] tokens within the prompt, while the “EOS” (End Of Sentence) column specifies the chosen terminal punctuation of the template. For example, a configuration with [MASK] set to 2 and EOS as ‘!’ results in the prompt ‘This sentence: “[X]” means [MASK][MASK] !’. In addition, the “Aggregation” column clarifies how the output vectors from the [MASK] tokens are combined, with “average” indicating an average pooling and “concatenate” signifying a sequential concatenation of these vectors.
The analysis reveals that omitting a terminal character in the manual template, BERT achieves a mere Spearman correlation coefficient score of 27.21 on the STS-B development set. In contrast, the introduction of a period, exclamation mark, or question mark substantially bolsters the representational capacity of BERT’s raw embeddings, elevating its score to reach over 70. Interestingly, the period emerges as the least effective among these punctuation marks. Additionally, the findings indicates that utilizing a single [MASK] token is suboptimal when the sentence concludes with a period or question mark. In other words, the EOL hypothesis of summarizing the original text into one word does not always hold true for discriminative models like BERT. Moreover, for templates featuring multiple [MASK] tokens, an average pooling of the output vectors consistently outperforms direct concatenation.
@@ 3.2 EOL in the Context of Generative Models
In Table 3, the “Fine-tuning” column corresponds to the experimental results obtained by employing the NLI dataset as training corpus, coupled with the application of contrastive learning and QLoRA for fine-tuning. Conversely, the “Direct Inference” column represents the Spearman correlation scores derived from leveraging the raw embeddings of PLMs. According to the experimental outcomes, the performance disparities among different templates post parameter updates are actually quite minimal. It could even be said that PromptSUM performs better than PromptEOL when fine-tuned. However, within direct inference scenarios, PromptEOL demonstrates a clear advantage. For instance, employing Mistral, PromptSUM attains an average Spearman correlation of 85.83 following supervised fine-tuning, slightly higher than PromptEOL’s 85.50. Nonetheless, in direct inference, PromptEOL significantly outperforms PromptSUM by 11.41 points (73.32 vs. 61.91).
@@ 3.3 Pretended CoT and Knowledge Enhancement
The experimental findings detailed in Table 3 demonstrate that while EOL notably enhances the semantic expressiveness of generative PLMs, the direct inference capabilities of 7B-scale models remain inferior to those of unsupervised fine-tuned SimCSE-BERT, which achieves an average Spearman correlation score of 76.25 across seven STS benchmarks. This score markedly surpasses the 70.03 performance metric of LLaMA2-EOL.
For further improvement, PromptEOL suggests adopting a one-shot format to provide the model with a clear example, so as to facilitate the text summarization process implied by EOL more effectively. While this method adeptly exploits auto-regressive models’ contextual learning capacity, the method’s reliance on specific examples limits its universality across PLMs of diverse types and sizes. Each example within the demonstration set must undergo evaluation against a validation set to identify the most suitable choice, a procedure that is resource-intensive and time-consuming. Alternatively, MetaEOL introduces a strategy involving the definition of meta-tasks, with ChatGPT generating corresponding templates for each. These templates are then amalgamated with the original sentence and processed through the PLM to derive multiple embeddings, thereby enabling the model to analyze various semantic aspects. This method, however, requires repeated encoding of sentences, diminishing efficiency.
This study aims to refine the quality of generative PLM raw embeddings in a more streamline and concise manner. To this end, we propose two plug-and-play prompting engineering techniques: Pretended CoT and Knowledge Enhancement. Their integration with EOL is illustrated in Table 4.
@@
Pretended CoT draws inspiration from Zero-shot CoT, which guides dialogue models to sequentially solve complex reasoning problems by appending “Let’s think step by step.” at the end of the prompt. Empirical evidence suggests that the this strategy effectively unlocks the potential of large-scale PLMs, enabling them to complete target tasks with higher accuracy. In pursuit of computational resources and efficiency, we chose not to use as large generative models as Zero-shot CoT. In fact, the majority of experiments in this paper are conducted on PLMs at or below 7B scale. Although these models possess some text continuation capabilities, the quality of the generated content is generally low. Therefore, rather than eliciting intermediate reasoning steps, we encourage a meticulous approach to sentence representation by prefacing EOL with “After thinking step by step,”.
@@
It is important to highlight that, to minimize operational complexities, the implementations of Pretended CoT and Knowledge Enhancement for all PLMs, regardless of their specific types and sizes, are completely consistent in our research. In forthcoming sections, we will demonstrate through specific experimental results that this unified configuration is sufficient to provide a stable performance improvement for the models.

### v2

@@ Abstract
The advent of generative Pre-trained Language Models (PLMs), exemplified by GPT [1], OPT [2], and LLaMA [3], marks a significant milestone in the field of Natural Language Processing (NLP). Boasting immense parameter sizes and extensive pre-training corpora, these models have demonstrated unparalleled proficiency in multi-task processing and zero-shot reasoning. Consequently, there is a growing interest in directly leveraging these PLMs for a wide array of downstream tasks, including the derivation of sentence embeddings.
@@
Over the past few years, research in this direction has predominantly revolved around discriminative models like BERT [4] and RoBERTa [5]. Nonetheless, the semantic space anisotropy inherent in these models [6] often mandates fine-tuning with considerable datasets to produce high-quality sentence embeddings. Among various fine-tuning approaches, contrastive learning, which aims to bring similar samples closer while distancing disparate ones, has proven to be the most effective. However, to prevent model collapse and ensure sufficient reference items for sentence vector calculation, contrastive learning methods necessitate large training batch sizes, thereby consuming massive computational resources.
@@
In response to this challenge, some cutting-edge efforts have shifted their focus to the recently developed generative models, hoping to leverage their advanced text comprehension abilities to directly encode input sentences without additional backward propagation. Given the disparity between sentence representation and auto-regressive language modeling, this endeavor demands sophisticatedly designed prompts. The pioneering work, PromptEOL [7], find that by introducing an Explicit One word Limitation (EOL) into the prompt and constructing a template akin to ‘This sentence: “[X]” means in one word: “’, it can significantly enhance the embedding obtained from generative models. Subsequent studies [8] [9] have followed this practice while also combining EOL with contrastive learning for fine-tuning Large Language Models (LLMs). Nevertheless, even with efficient fine-tuning strategies like QLoRA [10], training a neural network with a 7B-scale LLM as the backbone still consumes more GPU memory than fully fine-tuning a 110M-scale BERT. Moreover, we elucidate in subsequent sections that EOL chiefly benefits raw embeddings from generative models and is not requisite for discriminative models or fine-tuning generative models.
Therefore, to reconcile the demand for high-quality sentence embeddings with the need for computational resource conservation, it is advisable to adopt PLMs with relatively high parameter scale while avoiding gradient updates. Given that the quality of embeddings directly encoded from input sentences is generally low [7], this paper designs two simple yet effective prompt engineering methods: Pretended Chain of Thought (CoT) and Knowledge Enhancement, to ameliorate the sentence representation capabilities of generative models under direct inference settings. Both techniques involve appending a fixed prefix before the EOL to harness PLMs’ contextual learning ability to its fullest potential.
@@
Specifically, Pretended CoT, inspired by the design of Zero-shot CoT [12], mainly serves an “inspirational” role since we do not need the model to output intermediate reasoning steps. Knowledge Enhancement, on the other hand, provides explicit guidance to the model by conveying human experience in text summarization through prompts. Our comprehensive experimental outcomes affirm that Pretended CoT and Knowledge Enhancement not only optimize memory efficiency but also significantly elevate the performance of raw embeddings from generative models across multiple Semantic Textual Similarity (STS) benchmarks.
@@
Our investigations reveal that the widely applied Explicit One-word Limitation (EOL) is actually a sentence embedding derivation technique for direct inference scenarios with generative PLMs. It does not always yield optimal performance in discriminative models or fine-tuning generative models. This finding opens new avenues for future research on manual template design.
This paper proposes two concise and potent prompt engineering methods, Pretended CoT and Knowledge Enhancement, that markedly enhance the quality of embeddings produced by generative models on top of the original prompt.
@@ 2 Related Work
The exploration of sentence embeddings with PLMs constitutes a vibrant area of research within Natural Language Processing and Computational Linguistics. This body of work diverges into three primary streams: supervised fine-tuning, unsupervised fine-tuning, and direct inference. Each paradigm offers unique insights while presenting distinct challenges, as delineated below.
@@
Supervised Fine-tuning: This strand of research leverages high-quality annotated corpora to enhance models’ encoding capabilities. A landmark in this domain, Sentence-BERT [13] designs a Siamese neural network architecture with shared parameters and updates model via the NLI dataset [14, 15]. Following this, SimCSE [16] filters the NLI dataset and organizes it into triplets to adjust the sentence vector distribution of BERT through contrastive learning. Further studies like PromptEOL [7], AnglE [8], and DeeLM [9] extend this methodology to generative PLMs such as OPT [2], LLaMA [3], and LLaMA2 [17]. While supervised learning methods consistently outperform across a spectrum of benchmarks, their dependency on labeled datasets limits their applicability in a broader context.
@@
Unsupervised Fine-tuning: Focusing on the potential of models trained with unlabelled corpora, unsupervised fine-tuning has emerged as a fertile ground for innovation. This realm is particularly enriched by contrastive learning methods and data augmentation techniques. ESimCSE [18], for example, advances SimCSE by introducing random word repetition in the input text to construct positives for original sentences. PromptBERT [19], in contrast, utilizes different manual templates to create positive instances. Additionally, SSCL [20] demonstrates the efficacy of using intermediate layer outputs from PLMs as hard negatives in the InfoNCE Loss [21] framework. CoT-BERT [22], representing the current state-of-the-art, nearly matches the performance of supervised approaches across multiple STS benchmarks by relying solely on the intrinsic capabilities of PLMs. Despite these encouraging advancements, unsupervised fine-tuning methods do not inherently present advantages in memory efficiency, given their reliance on large batch sizes to facilitate effective contrastive learning.
@@
Direct Inference: Efficient sentence representation methods that directly utilize raw embeddings of PLMs without parameter updates are increasingly attracting attention in the NLP community. Among these, PromptBERT [19] showcases that meticulously designed templates tailored to mimic the Masked Language Modeling (MLM) task can exploit BERT’s original output vectors for commendable performance. Building on this insight, PromptEOL [7] targets generative PLMs and advocates explicitly prompting the model to condense the original sentence into a single word. This strategy markedly taps into the latent potential of generative models and surpasses the direct inference results of PromptBERT. Considering the multifaceted nature of refining textual semantics, MetaEOL [11] employs ChatGPT-4 to craft diverse templates, feeding them into the PLM and extracting the sentence representation through average pooling. Although the rationale behind this method is theoretically sound, encoding the same sentence multiple times with distinct templates diverges from the original intention of offline computation.
@@ 3 Methodology
In this chapter, we first demonstrate through pilot experiments in Sections 3.1 and 3.2 that the prevailing paradigm in current sentence representation research, Explicit One-word Limitation (EOL), primarily functions to enhance the direct inference effects of generative PLMs. It does not consistently yield superior outcomes in discriminative models or when fine-tuning generative PLMs. Therefore, explorations surrounding EOL should focus more on scenarios that directly apply PLM raw embeddings for inference. Subsequently, in Section 3.3, we provide detailed descriptions of two innovative prompt engineering techniques proposed in this paper, Pretended Chain of Thought (CoT) and Knowledge Enhancement, for further elevating the semantic expressiveness of PLMs’ raw embeddings.
@@ 3.1 EOL in the Context of Discriminative Models
The design philosophy of EOL stems from an analysis of inference outcomes when employing an unmodified BERT model on the STS-B development set [32]. This analysis contrasts two highly similar manual templates, distinguished solely by the presence or absence of a terminal period:
@@
Expanding upon this premise, our study embarks on a more granular investigation. We commence with a series of comparative experiments centered on BERT, examining the impact of the number of [MASK] tokens and the selection of terminal characters in templates. Adopting the identical experimental setup as PromptEOL, we utilize ‘This sentence: “[X]” means [MASK].’ as our baseline template and initialize BERT from its original pre-trained state with its performance measured on the STS-B development set via Spearman correlation coefficient. The corresponding results are presented in Table 1.
| [MASK] | EOS | STS-B dev |
| 1 | None | 27.21 |
| 1 | [SEP] | 32.54 |
| 1 | . | 73.61 |
| 2 | . | 73.94 |
| 3 | . | 74.50 |
| 4 | . | 71.83 |
| 1 | 77.59 | |
| 2 | 76.55 | |
| 3 | 75.99 | |
| 4 | 73.45 | |
| 1 | 75.32 | |
| 2 | 75.19 | |
| 3 | 75.37 | |
| 4 | 73.62 |
In Table 1, the column labeled “[MASK]” denotes the number of [MASK] tokens within the prompt, while the “EOS” (End of Sentence) column specifies the selected terminal punctuation of the template. For example, a configuration with [MASK] set to 2 and EOS as ‘!’ results in the prompt ‘This sentence: “[X]” means [MASK][MASK] !’. In cases where multiple [MASK] tokens are employed, we calculate the mean of their output vectors as the corresponding sentence representation.
The analysis reveals that omitting a terminal character in the manual template, BERT achieves a mere Spearman correlation coefficient score of 27.21 on the STS-B development set. In contrast, the introduction of a period, exclamation mark, or question mark substantially bolsters the representational capacity of BERT’s raw embeddings, elevating its score to reach over 70. Interestingly, the period emerges as the least effective among these punctuation marks. Additionally, the findings indicate that utilizing a single [MASK] token is suboptimal when the sentence concludes with a period or question mark. In other words, the EOL hypothesis of summarizing the original text into one word does not always hold true for discriminative models like BERT.
@@ 3.2 EOL in the Context of Generative Models
In Table 2, the “Fine-tuning” column corresponds to the experimental results obtained by employing the NLI dataset as training corpus, coupled with the application of contrastive learning and QLoRA for fine-tuning. Conversely, the “Direct Inference” column represents the Spearman correlation scores derived from leveraging the raw embeddings of PLMs. According to the experimental outcomes, the performance disparities among different templates post parameter updates are actually quite minimal. It could even be said that PromptSUM performs better than PromptEOL when fine-tuned. However, within direct inference scenarios, PromptEOL demonstrates a clear advantage. For instance, employing Mistral7b, PromptSUM attains an average Spearman correlation of 85.83 following supervised fine-tuning, slightly higher than PromptEOL’s 85.50. Nonetheless, in direct inference, PromptEOL significantly outperforms PromptSUM by 11.41 points (73.32 vs. 61.91).
@@ 3.3 Pretended CoT and Knowledge Enhancement
The experimental findings detailed in Table 2 demonstrate that while EOL notably enhances the semantic expressiveness of generative PLMs, the direct inference capabilities of 7B-scale models remain inferior to those of unsupervised fine-tuned SimCSE-BERT, which achieves an average Spearman correlation score of 76.25 across seven STS benchmarks [16]. This score markedly surpasses the 70.03 performance metric of LLaMA27b-EOL.
For further improvement, PromptEOL suggests adopting a one-shot format to provide the model with a clear example, so as to facilitate the text summarization process implied by EOL more effectively. While this method adeptly exploits auto-regressive models’ contextual learning capacity, its reliance on specific examples limits its universality across PLMs of diverse types and sizes. Each example within the demonstration set must undergo evaluation against a validation set to identify the most suitable choice, a procedure that is both resource-intensive and time-consuming. Alternatively, MetaEOL introduces a strategy involving the definition of meta-tasks, with ChatGPT generating corresponding templates for each. These templates are then amalgamated with the original sentence and processed through the PLM to derive multiple embeddings, thereby enabling the model to analyze various semantic aspects. This method, however, requires repeated encoding of sentences, diminishing efficiency.
This study aims to refine the quality of generative PLM raw embeddings in a more streamline and concise manner. To this end, we propose two plug-and-play prompt engineering techniques: Pretended CoT and Knowledge Enhancement. Their integration with EOL is illustrated in Table 3.
@@
Pretended CoT draws inspiration from Zero-shot CoT, which guides dialogue models to sequentially solve complex reasoning problems by appending “Let’s think step by step.” at the end of the prompt. Empirical evidence suggests that this strategy effectively unlocks the potential of large-scale PLMs, enabling them to complete target tasks with higher accuracy. In pursuit of computational resources and efficiency, we choose not to use as large generative models as Zero-shot CoT. In fact, the majority of experiments in this paper are conducted on PLMs at or below 7B scale. Although these models possess some text continuation capabilities, the quality of the generated content is generally low. Therefore, rather than eliciting intermediate reasoning steps, we encourage a meticulous approach to sentence representation by prefacing EOL with “After thinking step by step,”.
@@
It is important to highlight that, to minimize operational complexities, the implementations of Pretended CoT and Knowledge Enhancement for all PLMs, regardless of their specific types and sizes, are completely consistent in our research. In forthcoming sections, we will demonstrate through specific experimental results that this unified configuration is sufficient to provide a stable performance improvement for models.